OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 one fourth -one fourth -4

OpenStudy (ellecullen):

slope = x2-x1/y2-y1 formula so -3+5 / 7+1 2/8 = 1/4 the answer is one fourth Elle, High school senior

OpenStudy (mathstudent55):

Subtract -3 from -5 Subtract 7 from -1 Divide the first subtraction by the second subtraction